A population is estimated to have a standard deviation of 12. We want to estimate the population mean within 2 with a 99 percent level of confidence. How large a sample is required?
Answer #1

Hello Sir/ Mam

We want to estimate population mean within 2. Hence, clearly, Margin of error = 1

Hence,

12 1= 2* = 2.576 * vn

n = 2.576 * 12

n = (2.576 * 12)2 = 955.43 956

I hope this solves your doubt.

Feel free to comment if you still have any query or need something else. I'll help asap.

Do give a thumbs up if you find this helpful.